Ok, My son came home from school today with Math homework. He is in Algebra 1 and they have requested he have a TI-83. I intend to buy this for him but I haven't yet and he has homework tonight. The teacher only explained how to do the problem on the calulator not on paper. I know it can be done on paper but I really do not remember how to do it. I looked in his math book and the book only goes to Chapter 12 and the worksheet is for Chapter 13. I can't figure out how to solve these problems and I'm hoping for some help. The question reads as follows: Find the solution set for each equation, given the replacement set.
One of the problems(not one he has to do) read as follows:
y=4x=1; {(2,-1), (1,5), (9,2), (0,1)}.
If you could explain how to do this problem I'm sure we can do the problems he has to turn in for homework.
Thank you.


Answer by stanbon(57214)

Find the solution set for each equation, given the replacement set.
One of the problems(not one he has to do) read as follows:
y=4x+1; {(2,-1), (1,5), (9,2), (0,1)}.
--------------
Each of the pairs gives you values for x and for y .
You are to take one pair at a time and insert them into the equation to
see if you get a true result, as follows:
Looking at (2,-1) you would get:
-1=4*2+1
-1=8+1
That is wrong
-------------------
Looking at (1,5) you would get:
5=4*1+1
5=4+1
5=5
That is right so that pair is a solution for the equation
------------------------------------------
Looking at (9,2) you would get:
2=4*9+1
2=36+1
2=37
Not a solution for the equation.
----------------
Try the other pairs and see what you get; good luck.
======================
cheers,
Stan H.